Determine whether the given series are absolutely convergent, conditionally convergent or divergent. (same answers can be used multiple times) Determine whether the given series are absolutely convergent, conditionally convergent or divergent. (-1)(2n +3n2) 2n2-n is
Answer #1

(-1)(2n +312) 2n²_n nel an= (2n+3 n3) 2narni anti (-1)+1 ( 2 6n +1) + 3 (n+1)2] 2 (n+1) ² (n+1) (-199+ [2(0+1) +3 (n+1)?) 2hini (30+5)(1-21) (217+1 )(3n+2) о - и (uz-1)(stus) zu hui (2n+1)(3n+2) 72 3nts n -) (120) u hem no (20+) ( 3n+2 u 3+5lm - Lu(3) Z (-1)(n +2) 2n²1 n=1 On = (-1) (+2) 2n2-) L = lim anti an lim (-1)(n+1+2) 2(n+1)-1 (-1) (n+2) 2n2-1 n+1 = lim = (-1)(- (n+3) (2n2-1) - lim n2 no 2(n+1) 21 = lim И со 2 n+1 1 + 2 h2 -) (42) -(+ ) (2-6) n) lim - (+2)(2-6) no ( 2 (1+4)*72) (+) -